OpenStudy (anonymous):

The equation sin((pi/3)-x) is equal to: a. 1/2(√3cosx-sinx) b. 1/2(√3cosx+sinx) c. 1/2(cosx+√3sinx) d. 1/2(cosx-√3sinx)

OpenStudy (zepp):

Use the difference formula to solve :]

OpenStudy (zepp):

\[\large \sin(\theta \pm\phi)=\sin \theta \cos \phi\pm\cos\theta\sin\phi\]